The problem is this:
2. As you are leaving the community, you notice another new community just down the street. Because you are in the area, you decide to inquire about it.
a)The sales representative here tells you they also have two floor plans available, but they only have 38 homes available. Write an equation that illustrates the situation. Use x and y to denote floor plan #1 and floor plan #2 respectively. (( X + y = 38) is what i thought the equation would be)
b) The representative tells you that floor plan #1 sells for $175,000 and floor plan #2 sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in part a.
((175,000x +200000y =72,000,000)is what i put for the equation)
c) Use elimination to determine how many houses with each floor plan are available. Explain how you arrived at your answer.
I have done this maybe 10 times. i keep coming up with x = 2576 but that does not make sense when i put that back into the orginal equation. Please Help! I have till 3 am est to turn it in. Thanks!

Answer by Mathtut(3670)

2a)x + y = 38
=======

2b) We have x houses @ $175,000 and y houses @ $200,000, with a total worth of $7,200,000.

The equation representing this relationship is,

175,000x + 200,000y = 7,200,000

simplifying the numbers gives us,

175x + 200y = 7,200
7x + 8y = 288

Our system of simultaneous equatiuons is,

7x + 8y = 288
x + y = 38

Multiply the 2nd equation by 7,

7x + 8y = 288
7x + 7y = 266

eliminate the x by subtracting the 2nd eqn from the 1st.

y = 22
=====
x + 22 = 38 (using y = 22)
x = 16